OpenStudy (becca9898):

What are the factors of f(x)? f(x) = x^3 + 5x^2 - x - 5

OpenStudy (agent47):

\[f(x) = x^3+5x^2-x-5=x^2(x+5)-x-5=x^2(x+5)-(x+5)=(x+5)(x^2-1)\]

OpenStudy (agent47):

\[f(x) = x^2(x+5)-(x+5)=(x+5)(x^2-1)=(x+5)(x-1)(x+1)\]

OpenStudy (becca9898):

I don't understand..

OpenStudy (agent47):

What don't you understand?